Find the measure of an angle that is 30 degrees less than its suppplement
Answer by algebrahouse.com(1659)

x = the angle
180 - x = its supplement {supplementary angles add up to 180°}

x = 180 - x - 30 {angle is 30° less than its supplement}
x = 150 - x {combined like terms}
2x = 150 {added x to each side}
x = 75° {divided each side by 2}

For more help from me, visit: www.algebrahouse.com